An element with mass 780 grams decays by 17.1% per minute. How much of the element is remaining after 19 minutes, to the nearest 10th of a gram?



1
Expert's answer
2021-03-08T15:28:42-0500

element remaining"=mass \times(1-r/100)^t=780\times(1-17.1/100)^{19}=780\times 0.829^{19}=22.1g"